OpenStudy (abbycross167):

Write the slope-intercept (y = mx + b) form of the equation from the information given. The line goes through the points (-1, -7) and (9, -2) So I need help... can someone try to show me how to do this?

OpenStudy (smilie2015):

first you have to find slope through formula m=\frac{{y2 - y1}}{{x2 - x1}} here x1= -1,y1=-7,x2=9,y2=-2 put the values in equation

OpenStudy (smilie2015):

\[\frac{{y2 - y1}}{{x2 - x1}}\]

OpenStudy (smilie2015):

you will get 5/10 =1/2 so slope is 1/2 now y-y1=m(x-x1) y+7=1/2(x+1) y+7=1/2x+1/2 y=1/2x+1/2-7

OpenStudy (smilie2015):

\[y = \frac{1}{2}x - \frac{{13}}{2}\]

OpenStudy (abbycross167):

so that's the equation ?

OpenStudy (smilie2015):

yup

OpenStudy (abbycross167):

oh ok, thank you!!! soo much! haha you make me smile :) (username pun)

OpenStudy (smilie2015):

hahaha u r welcome :)